# Dynamic phase transition into a mixed-CDW state in 1T-TaS2 via a thermal quench

### v2

@@ II Results
Scanning tunneling microscopy (STM), measurements confirm the multi-domain nature of our quenched samples. Figure 3 a shows a characteristic $80\times 80$ nm2 area in a quenched sample. Each domain shows the atomic reconstruction of the sulfur layer due to low-temperature CDW, as shown in Fig. 3 c with $V\!=\!380$ mV, $I\!=\!0.35$ nA, scale bar length $\!=\!2$ nm). We note that the in-plane momentum resolution of our STM experiments prevents us from distinguishing between the C-CDW and the H-CDW-like domains. However, from the Fourier transform of STM images with atomic resolution, we extract a rotated satellite peak $\phi\!=\!12.8(5)^{\circ}$, that reflects the presence of domains of opposite chirality (See extended data).
@@
Electronic structure of the bulk quenched samples. We now bring attention to the low-temperature electronic structure of the mixed-CDW phase. While the origin of the insulating nature of the C-CDW remains controversial Petocchi et al. (2022); Wang et al. (2020), it is well established that the c-axis stacking of CDW domains strongly influences the electronic structure of $1T$-TaS2 Ritschel et al. (2018); Lee et al. (2019); Butler et al. (2020); Ritschel et al. (2015). Moreover, recent STM and scanning tunneling spectroscopy (STS) studies of the stack shift between in-plane domains in consecutive layers have identified interlayer coupling as the key ingredient that determines the band structure of the low-temperature CDW phase in $1T$-TaS2 Cho et al. (2017); Ma et al. (2016); Park et al. (2021); Gerasimenko et al. (2019); Park et al. (2023). Thus, we use STS to study the low-temperature electronic state of the quenched samples. $dI/dV$ curves taken at $T\!=\!10$ K within six different domains, Fig. 3 a, reveal a reduction of the $E_{g}\!=\!300$ meV gap of the C-CDW Campbell et al. (2024) towards a state with a reduced but non-zero density of states at the Fermi level (Fig. 3 c). While a transition to a semimetallic state is consistent with that observed for the H-CDW Ma et al. (2016); Gerasimenko et al. (2019), the density of states maintains the two pronounced peaks at around $V_{B}=-300$ meV and $V_{B}=150$ meV, that have been associated with the lower and upper Hubbard bands Perfetti et al. (2005); Kim et al. (1994); Cho et al. (2015). Despite the finite in-plane density of states at $E_{F}$ and the formation of metallic domain walls (Fig. 3 b, curve 3), electrical transport in the mixed-CDW sample resembles that of the as-grown crystals.
@@
Figure 4: Dynamic phase transition with two competing order parameters in $1T$-TaS2. a Schematic free energy landscape for a dynamical phase transition with two competing order parameters. A quench can stabilize a local minimum, different than the global minimum with order parameter $\psi_{1}$, corresponding to a secondary order parameter $\psi_{2}$. b-d Schematic of the possible low-temperature groundstates of $1T$-TaS2 as a function of quenching rate. c C-CDW, d mixed-CDW state and e H-CDW.
@@ III Discussion
In summary, our results demonstrate that a mixed-CDW phase characterized by C-CDW and H-CDW domains of opposite chirality can be stabilized by a thermal quench through $T_{\mathrm{I-NC}}$. Our STS data reveals a semimetallic in-plane electronic density of states, while out-of-plane CDW stacking disorder prevents bulk metallic behavior.
@@
It is compelling to compare our observations with theories for first-order dynamical phase transitions with two competing order parameters Sun and Millis (2020), in which the order parameter of the C-CDW corresponds to the global minimum of the free energy landscape. Under this hypothesis, the H-CDW phase is a local minimum of the free-energy landscape as sketched in Fig. 4 a. This provides a natural explanation for the large C-CDW domains for slow quenches, as sketched in Fig. 4 b and c. Such a free energy landscape allows a mixed inhomogeneous phase to be accessed at intermediate but adiabatic quenching rates in which a single temperature can be defined for all the different degrees of freedom in the material. This state is characterized by the coexistence of trapped regions of the two competing orders Del Re et al. (2016) and enhanced lifetime due to the transiently created domain walls Sun and Millis (2020). Our data is consistent with such a scenario: in $1T$-TaS2 a thermal quench stabilizes a mixed phase of H-CDW and C-CDW order parameters up to $T\!=\!180$ K in cooling and $T\!=\!210$ K in warming (Fig. 4 d).
@@
We note that all our thermal quenching attempts were unsuccessful in engineering a sample without signatures of the C-CDW state, as sketched in Fig. 4 e. It is likely that the quench protocol we describe here results in a nonuniform quench rate that is locally too slow over some regions of the sample to allow for the nucleation of the H-CDw domains. Within this context, previous ultrafast stabilization of the H-phase in thin flakes Stojchevska et al. (2014); Stahl et al. (2020); Vaskivskyi et al. (2015); Hollander et al. (2015); Ma et al. (2016); Cho et al. (2016); Venturini et al. (2022); Tsen et al. (2015), can be understood as a non-thermal ultrafast quench Sun and Millis (2020). Our results are consistent with proposals in which the absorption of energy by the electronic degrees of freedom suppresses the C-CDW order parameter. The consequent relaxation of the non-thermal electrons via the lattice leads to an ionic displacement from that of the C-CDW state into that of the H-CDW Maklar et al. (2023) allowing for the correlations of the H-CDW order parameter to grow Sun and Millis (2020). Thus, our results suggest that the minimum of the free energy landscape associated with the H-CDW order parameter already exists at the thermodynamic equilibrium of $1T$-TaS2 and can be accessed with a sufficiently fast adiabatic quench (Figure 3 a).
@@ V Acknowledgements
A.d.l.T acknowledges helpful conversation with Michael Buchold, Yue Cao, Martin Claassen, Greg Fiete, Simon Gerber, and Dante M. Kennes. Work performed at Brown University by A.d.l.T., Q.W. and K.W.P. was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-SC0021223. P.M.V. and D.B. acknowledge support from NSF DMR 2226097 and the Mason Graduate Division’s Presidential Scholarship Program”. This work is based on research conducted at the Center for High-Energy X-ray Sciences (CHEXS), which is supported by the National Science Foundation (BIO, ENG and MPS Directorates) under award DMR-1829070.
@@
A.d.l.T and Q. W. synthesized and characterized the single crystal samples. A.d.l.T., Q. W., K. W. P. and J.P.C.R. performed the High Dynamic Range x-ray scattering measurements. All x-ray experimental data was analyzed and interpreted by A.d.l.T and K .W. P.. B. C., J. V. R., and S. M. H. performed and analyzed the STM and STS measurements. D. B. and P.M.V. performed the cryo-Raman experiments.‘ A.d.l.T. and. K. W. P wrote the manuscript with input from all co-authors.

### v3

@@ II Results
Scanning tunneling microscopy (STM), measurements confirm the multi-domain nature of our quenched samples. Figure 3 a shows a characteristic $80\times 80$ nm2 area in a quenched sample. Each domain shows the atomic reconstruction of the sulfur layer due to a low-temperature CDW, as shown in Fig. 3 c with $V\!=\!380$ mV, $I\!=\!0.35$ nA, scale bar length $\!=\!2$ nm). We note that the in-plane momentum resolution of our STM experiments prevents us from distinguishing between the C-CDW and the H-CDW-like domains. However, from the Fourier transform of STM images with atomic resolution, we extract a rotated satellite peak $\phi\!=\!12.8(5)^{\circ}$, that reflects the presence of domains of opposite chirality (related by a mirror operation along one of the high symmetry crystallographic directions) (See extended data). Our STM data resembles the in-plane inhomogeneity observed in optically and electrically driven $1T$-TaS2, where multiple domains of opposite phase have been reported Gerasimenko et al. (2019); Ma et al. (2016); Maklar et al. (2023). We note that the smaller domain density prevented us from performing a comprehensive in-plane phase configuration analysis of the chiral domains, but we cannot exclude the presence of phase domains (related by a translation operation) Gerasimenko et al. (2019); Ma et al. (2016); Maklar et al. (2023).
@@
Electronic structure of the bulk quenched samples. We now bring attention to the low-temperature electronic structure of the mixed-CDW phase. While the origin of the insulating nature of the C-CDW remains controversial Petocchi et al. (2022); Wang et al. (2020), it is well established that the c-axis stacking of CDW domains strongly influences the electronic structure of $1T$-TaS2 Ritschel et al. (2018); Lee et al. (2019); Butler et al. (2020); Ritschel et al. (2015). Moreover, recent STM and scanning tunneling spectroscopy (STS) studies of the stacking shift between in-plane domains in consecutive layers have identified interlayer coupling as the key ingredient that determines the band structure of the low-temperature CDW phase in $1T$-TaS2 Cho et al. (2017); Ma et al. (2016); Park et al. (2021); Gerasimenko et al. (2019); Park et al. (2023). Thus, we use STS to study the low-temperature electronic state of the quenched samples. $dI/dV$ curves taken at $T\!=\!10$ K within six different domains, Fig. 3 a, reveal a reduction of the $E_{g}\!=\!300$ meV gap of the C-CDW Campbell et al. (2024) towards a state with a reduced but non-zero density of states at the Fermi level (Fig. 3 c). While a transition to a semimetallic state is consistent with that observed for the H-CDW Ma et al. (2016); Gerasimenko et al. (2019), the density of states maintains two pronounced peaks at around $V_{B}=-300$ meV and $V_{B}=150$ meV, that have been associated with the lower and upper Hubbard bands Perfetti et al. (2005); Kim et al. (1994); Cho et al. (2015). Despite the finite in-plane density of states at $E_{F}$ and the formation of metallic domain walls (Fig. 3 b, curve 3), electrical transport in the mixed-CDW sample resembles that of the as-grown crystals.
Figure 4: Dynamic phase transition with two competing order parameters in $1T$-TaS2. a Cooling rate estimates for an optically induced quench (orange) and solutions of the heat diffusion equation for the slow (red) and fast (blue) cooling schemes used in this work (See Methods and Extended data). b Meanfield-only free energy functional for the two competing CDW orders, with a competition factor $d_{0}=0.8$, corresponding to an adiabatic quench. c 2D projection of the energy functional in b. d Meanfield-only free energy functional for the two competing CDW orders, with a competition factor $d_{0}=6.1$, corresponding to laser quench. Fast quenches correspond to larger values of $d_{0}$. e 2D projection of the energy functional in d. The yellow and red markers indicate local and global minima, respectively. f Computed phase diagram as a function of the competition factor, $d_{0}$, and mass term, $a_{0}$. The mixed CDW phase can only be stabilized in a narrow range of intermediate-rate quenches. Faster quenches (optical or electrical pulses) lead to the suppression of the C-CDW phase and the emergence of the H-CDW phase
@@ III Discussion
In summary, our results demonstrate that a mixed-CDW phase characterized by C-CDW and H-CDW domains of opposite chirality can be stabilized by a thermal quench through $T_{\mathrm{I-NC}}$. The required thermal quench is at an intermediate rate ($\approx 120$ K/s) based on heat diffusion calculations (see Methods), faster than the initial quench that $1T$-TaS2 crystals experience during growth ($\leq 2$ K/s) but slower than the changes to the electronic temperature induced by a laser pulse ($\geq 1$ K/fs) Zong et al. (2018); Stojchevska et al. (2014) (Figure 4) a. We note that the chiral splitting (relative rotation) of the two order parameters enables us to clearly resolve the mixed-CDW phase. If the H-CDW and C-CDW phases exhibited the same chirality, the small difference in their ordering wavevectors would lead to overlapping satellite peaks, complicating the observation of this novel state. Our STS data reveals that this mixed state hosts a semimetallic in-plane electronic density of states, while out-of-plane CDW stacking disorder prevents bulk metallic behavior.
@@
It is compelling to compare our observations with theories for dynamical phase transitions Sun and Millis (2020); Sefidkhani et al. (2024) and how domain formation influences the competition between coexisting order parameters. Thus, to describe the low temperature state after a quench, we consider a mean-field energy functional McMillan (1975) within the Landau-Ginzburg theory including, among others (see the Supplementary Information for more details of the calculations), a $d_{0}\bar{\phi}^{2}_{C}\bar{\phi}^{2}_{H}$ term in which $d_{0}$ parameterized the competition between the C-CDW and the H-CDW phase. The order parameter is the spatial charge modulation, $\alpha(\vec{r})=\Re(\phi_{C}e^{i\vec{Q}_{C}\cdot\vec{r}}+\phi_{H}e^{i\vec{Q}_{H}\cdot\vec{r}})$ in which the amplitude of the order parameter for the C-CDW, ${\phi}_{C}$, corresponds to the global minimum of the free energy landscape for slow cooling protocols, and $\phi_{H}$ is the amplitude of the H-CDW order parameter. We note that our model does not include the order parameter for the NC-CDW as the initial quench from $T_{\mathrm{I-NC}}$ down to low temperature effectively skips over the intermediate CDW state. Morevoer, the H-CDW phase can only be stabilized from the C-CDW phase Stojchevska et al. (2014) indicating that the NC-CDW is not necessary to discuss the competition between these two states. Such free energy landscape allows a mixed inhomogeneous phase to be accessed at intermediate but adiabatic quenching rates in which a single temperature can be defined for all the different degrees of freedom in the material. This state is characterized by the coexistence of trapped regions of the two competing orders Del Re et al. (2016) and enhanced lifetime due to the transiently created domain walls Sun and Millis (2020).
[…]
@@
Our data is consistent with a scenario in $1T$-TaS2 in which a thermal quench stabilizes a mixed phase of H-CDW and C-CDW order parameters up to $T\!=\!180$ K in cooling and $T\!=\!210$ K in warming. We note that coexistence of the H-CDW and C-CDW phase in ultrafast switched samples has also been recently reported Burri et al. (2024). Moreover, all our thermal quenching attempts were unsuccessful in engineering a sample without signatures of the C-CDW state indicating that sufficiently large values of $d_{0}$ where not achieved. It is likely that the quench protocol we describe here results in a nonuniform quench rate that is locally too slow over some regions of the sample to allow for the nucleation of the H-CDW domains. Within this context, previous ultrafast stabilization of the H-phase in thin flakes Stojchevska et al. (2014); Stahl et al. (2020); Vaskivskyi et al. (2015); Hollander et al. (2015); Ma et al. (2016); Cho et al. (2016); Venturini et al. (2022); Tsen et al. (2015), can be understood as a non-thermal ultrafast quench Sun and Millis (2020); Sefidkhani et al. (2024) leading to high domain densities and large values of $d_{0}$. Our results are consistent with proposals in which the absorption of energy by the electronic degrees of freedom suppresses the C-CDW order parameter. The consequent relaxation of the non-thermal electrons via the lattice leads to an ionic displacement from that of the C-CDW state into that of the H-CDW Maklar et al. (2023) allowing for the correlations of the H-CDW order parameter to grow Sun and Millis (2020). Thus, our results and Landau-Ginzburg theory calculations suggest that the minimum of the free energy landscape associated with the H-CDW order parameter already exists at the thermodynamic equilibrium of $1T$-TaS2 and can be accessed with a sufficiently fast adiabatic quench (Figure 4 a).
@@ V Acknowledgements
A.d.l.T acknowledges helpful conversation with Michael Buchold, Yue Cao, Martin Claassen, Simon Gerber, and Dante M. Kennes. Work performed at Brown University by A.d.l.T., Q.W. and K.W.P. was supported by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, under Award Number DE-SC0021223. P.M.V., S.M.H., and D.B. acknowledge support from NSF DMR 2226097 and the Mason Graduate Division’s Presidential Scholarship Program. This work is based on research conducted at the Center for High-Energy X-ray Sciences (CHEXS), which is supported by the National Science Foundation (BIO, ENG and MPS Directorates) under award DMR-1829070. G.A.F. and Y. M. gratefully acknowlege support from NSF DMR-2114825, and G.A.F. from the Alexander von Humboldt Foundation.
@@
A.d.l.T and Q. W. synthesized and characterized the single crystal samples. A.d.l.T., Q. W., K. W. P., and J.P.C.R. performed the High Dynamic Range x-ray scattering measurements. All x-ray experimental data was analyzed and interpreted by A.d.l.T and K .W. P.. B. C., J. V. R., and S. M. H. performed and analyzed the STM and STS measurements. D. B. and P.M.V. performed the cryo-Raman experiments. Y. M. and G.A.F. performed the Landau-Ginzburg free energy calculations. A.d.l.T. and. K. W. P wrote the manuscript with input from all co-authors.